# Language Models Surface the Unwritten Code of Science and Society

### v4

@@ 6 Discussions
It is important to emphasize that our goal is not to advocate for the automation of judgment tasks by LLMs. Rather, our primary goal is to present a position through a conceptual framework — namely, to exploit the biases and preference rationale embedded in LLMs by human discursive culture, then self-amplifying, reproducing, and explicitly identifying who might be harmed by these biases. Much of the existing research has focused on more explicit, easily identifiable forms of biases associated with social dimensions such as gender and race [47]. Researchers have developed methods to bypass the barriers when the model refuses to disclose its biases, such as using associated words [7]. However, they constitute the “prior” itself. These biases and heuristics still manifest superficially and can be uncovered through direct or indirect cues. In contrast, what we focus on in this paper is the "posterior" and how it diverges from the prior. Surprisingly, the results show that even when an LLM’s prior (i.e., its stereotypes) is "good", its posterior can still be "bad" - let alone when its prior is already "bad". When acting as reviewers, LLMs appear to draw on two distinct systems to explain differences in quality signals. One is more normative and less confident in the evaluation process. The other is more narrative-driven, context-sensitive, and ultimately more confident and consistent. When judgment can be achieved through both systems, LLMs are more likely to apply the second system across the board as heuristics because they more consistently explain quality differentials. Humans, like LLMs, may also favor contextual criteria in evaluation, but they articulate this shift in terms of normative standards.
@@
To what extent does an LLM’s shift from prior to posterior actually reflect human reasoning? Humans may never explicitly confess to using narrative or contextual heuristics in reviewing and other social dimensions, yet we have fairly strong evidence that they do. Prior survey-based studies show that reviewers in the social sciences place substantially greater emphasis on “interpretive challenges that question a study’s framing” than on methodological rigor [59, 56]. We push LLMs to articulate such emphasis and generalize this insight to conference papers in technical fields, where - owing to normative expectations around methodological rigor and SOTA performance - reviewers are less inclined to make such concerns explicit  [^4] https://hackingsemantics.xyz/2020/reviewing-models/ but use "lazy" words as a mask precisely in NLP [48]. Moreover, if human score variance could be understood as comprising both explicit and implicit components, with LLMs’ prior-to-posterior shift accounting for nearly all variance (as shown in Section 3 "iterative search") while human-stated explicit judgments map only to the prior, then the posterior can reasonably be taken to capture the implicit part.
@@
Another view holds that even if we acknowledge the existence of unwritten codes, making them explicit serves little purpose and may even enable people to “game the system.” But this view neglects the crucial balance between transparency and normativity. The fear that publicizing unwritten codes will be exploited implicitly assumes that it is preferable for such rules to continue operating in the shadows. In reality, the opposite is true: hidden rules primarily benefit those who already occupy advantaged positions or happen to be in the know. For example, researchers from elite institutions who are familiar with Western academic norms naturally understand how to write and frame a compelling “story” [13] - they have long benefited from tacit rules without needing them to be made explicit. The purpose of transparency is not to provide a cheat sheet, but to allow marginalized groups to understand the hidden barriers they face and thereby push institutions to reflect on and revise those standards. If people do game the system after unwritten codes are revealed, it is not a failure of transparency — it is evidence that those codes themselves need institutional correction.
@@
Can LLMs replace the peer review and other judge processes? Our goal is not to provide a definitive answer – although certainly not yet. Nevertheless, our work suggests that insofar as LLMs function as cultural technologies that encode societal norms [11, 16], then their "cognitive shift" in peer review reflects analogous unacknowledged shifts in human reviewers. That said, while LLM-generated reviews have been proven useful and cover broader and more specific aspects than human reviews [37, 69, 60] – likely including those humans may be unwilling or unable to disclose – LLM reviewers still lack constructive feedback essential for critique when judging a standalone paper. Pairing papers for comparative judgment of feature existence and magnitude in some sense addresses this concern, as demonstrated here and in previous experiments [7].
What should we do at the implementation level? LLMs themselves cannot become wholly unbiased. They can, however, present multiple perspectives, label potential biases, and even collaborate in a multi-model "ecosystem," thereby approximating a more neutral environment open to public scrutiny. In addition to making the unwritten codes spoken out by LLMs more explicit in rule-making, it is thus equally important to introduce diverse dimensions of perspectives. Taking science as an example, the perceived rhetorical value often depends on the reviewer’s proximity to the research [8]. We could therefore adopt a complementary–neutral–proximal reviewer composition. Through such collective calibration mechanisms [54], humans can not only mitigate but also heighten their awareness of potential biases, reconcile multiple cognitive perspectives, and maintain balance in group composition, thereby ensuring that different aspects of society benefit in turn.
@@
We used AI assistants to expedite the coding process. All code snippets produced by AI assistants were verified by the first author before they were incorporated. For writing, we only used AI assistants to check grammar. No external ethics review was required, as the experiments solely analyzed scientific papers, thereby posing no potential harm to individuals across any demographic or geographical group. Our study uses LLMs to surface implicit biases and unwritten evaluative codes in peer review. While this promotes transparency and fairness, it may also risk reinforcing stereotypes if misused. We mitigate this by anonymizing data, framing the method as diagnostic rather than prescriptive, and highlighting the importance of responsible use. All experiments were conducted using two NVIDIA 40-GB A100 GPUs.
@@ Appendix B Human annotation experiments
Here, we complement the labeling methods (LLM labeling (major) and the dictionary method (complementary)) with human annotation. Three master’s students specializing in social computing were involved to label the reviewer comments using exactly the same prompt as GPT-4o (if reviewers’ attitudes toward the aspects in a given hypothesis were positive = 1 or negative = –1, they were considered as “mentioning”, or 0 - not mentioned at all). The annotators are from Asia: two males and one female. They labeled independently, without using AI assistance and without communicating with one another. We find that one person shows high consistency with LLM labels, with an overlap of 0.76. Their majority-voted label across annotators shows an overlap with LLM labels of 0.7. These results show the reliability of LLMs in annotating text data.
@@ Appendix C Results using other LLMs are consistent
To increase the generalizability of our work, we explored other popular open-sourced or closed models including deepseek-chat-v3-0324  [^5] https://api-docs.deepseek.com/news/news250325 (the full set of experiments), LLaMA-3.3-70B-Instruct  [^6] https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, Claude-Sonnet-4  [^7] https://www.anthropic.com/news/claude-4, Claude 3 Haiku  [^8] https://www.anthropic.com/news/claude-3-haiku, Mistral Medium 3.1  [^9] https://openrouter.ai/mistralai/mistral-medium-3.1, and Gemini 2.0 Flash  [^10] https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash. Overall, we find our conclusions to be robust. Specifically:
Alignment with human comments: Deepseek also shows a relatively consistent set of priors as OpenAI models, and both align with human priors about the definition of “good science”, compared to the low alignment for posteriors VS human comments. Human reviewers tend to explicitly reward aspects that moderately align with LLMs’ normative priors: OpenAI correlation = 0.49; deepseek 0.28. By contrast, human reviewers generally avoid articulating posteriors in their review comments (OpenAI correlation = - 0.14; Deepseek correlation = –0.16), but both models do articulate these posteriors.
Consistent priors: we also find that models show relatively consistent priors compared to each other and humans (not only DeepSeek). Correlation: Llama-human = 0.49; Claude4-human 0.47, Mistral-human 0.41; Claude 3-human 0.41.
Shift from prior to posterior: During the transition from prior to posterior, we still observe a systematic shift from rigor-related aspects toward storytelling perspectives. All models demonstrate surprisingly high consistency in using certain heuristics such as “whether the paper tells a good interdisciplinary story”, and across Llama, Mistral, Deepseek, Gemini, and Claude models, they show on average 89% agreement with Open-AI’s judgment. This hypothesis also predictably gained the highest prior-to-posterior attention gain: interdisciplinary integration (OpenAI top-1; DeepSeek top-1), contextualization (DeepSeek top-2; OpenAI top-2), and readability (OpenAI top-4; DeepSeek top-4) are all systematically overestimated in the posterior belief shift. By contrast, rigor is consistently underestimated (OpenAI lowest top-1; DeepSeek lowest top-1), as is attention to bias in evaluation (OpenAI lowest top-2; DeepSeek lowest top-2), along with reproduction and implementation details (OpenAI lowest top-5; DeepSeek lowest top-3).
@@ Appendix E Fidelity experiments
Summarization using methods from [71] demonstrates high fidelity compared to the raw text. In Figure 7, panel (a), we use the embedding model from [55], specifically pretrained on scientific papers’ citation relations using contrastive learning. We embed both the summarizations and their corresponding raw texts in each section. To visualize and compare the distribution of these embedding areas, we reduce their dimensionality to two. We then estimate the boundary of each region using a 5% Gaussian kernel density estimate, drawing contours that enclose the densest 95% of data points in each embedding set. If the summarizations contain significant hallucinations, we would expect their embeddings to cover meaningfully different regions than those of the raw texts. However, in panel (a), we find that the distributions largely overlap, suggesting strong semantic alignment. In panel (b), we compute the cosine similarity between the two (raw, summarized) embeddings and find that most values are above 0.95.
@@
We conducted a complementary robustness check to assess whether LLMs make consistent judgments when directly reading full texts, compared to the extended summarization used in the paper for producing generalizable hypotheses under the limited context window. If using full texts from parsed PDFs, the results are essentially identical to those obtained with extended summaries (given their high fidelity), yielding a consistency score of 0.9 (i.e., the same judgment) across 1,000 judgments.
@@ Appendix G Related work
Hypothesis Generation: Recent studies have explored the capabilities of LLMs in generating novel, testable, and interpretable hypotheses across a range of domains. Zhou, Liu et al. [40, 39, 72] proposed a framework that prompts LLMs to generate hypotheses from labeled datasets, beginning with a few examples. These initial hypotheses are iteratively refined using a reward mechanism inspired by the upper confidence bound algorithm from multi-armed bandits, balancing exploration and exploitation. Other works have integrated external knowledge — such as scientific literature, knowledge graphs, and raw web corpus — into the hypothesis generation process [66, 28, 63, 64, 67], improving the accuracy and novelty of LLM-generated hypotheses while reducing hallucinations. This line of work demonstrates the potential of LLMs in advancing scientific discovery. Beyond computer science, other disciplines are also beginning to explore these possibilities. Ludwig and Mullainathan [42] highlighted the promise of LLMs as tools for hypothesis generation in economics, arguing that machine learning models can uncover patterns and relationships that traditional econometric methods may miss. Manning, Zhu, and Horton [43] presented a simulation-based approach for automatically generating and testing social science hypotheses in silico. In each simulated agent-interaction case, causal relationships are both proposed and tested by the LLM system, with results aligning well with those derived from economic theory, particularly in predicting the signs of estimated effects. In medicine, LLMs have been used to hypothesize synergistic drug combinations for breast cancer treatment, with several predictions later validated in laboratory experiments [10], showcasing the potential of LLMs to surface effective, previously overlooked therapeutic strategies.
@@
Implicit Bias in LLMs: Concerns about bias in AI systems have long focused on the ways these technologies can absorb and amplify harmful content from large-scale human-generated datasets – particularly prejudices related to race, gender, and other social dimensions [47]. Large language models (LLMs), trained on vast corpora of human text, are especially prone to perpetuating and even intensifying these biases [65, 18]. Much of the existing research on LLM bias has concentrated on explicit, easily identifiable forms of bias – such as derogatory language, disparate system performance, erasure, exclusionary norms, misrepresentation, stereotyping, and toxicity [18, 15]. Meanwhile, efforts in AI alignment, including outer alignment, inner alignment, and interpretability, have sought technological strategies for migrating such risks [52, 44, 29]. Despite progress in addressing explicit bias, unwritten codes in LLMs’ "cognitive" process remain understudied. Like human cognition, LLMs may exhibit subtle, unconscious forms of bias even when explicit prejudices are absent [52, 38]. These implicit biases are often harder to detect yet can have deeper and more pervasive effects on model outputs and downstream decision-making. Thus, systematically identifying these “unwritten codes" in LLMs is crucial to ensuring fairness, accuracy, and trustworthiness [38]. One method worth mentioning in this topic is [7]. The authors draw on experimental paradigms from the Implicit Association Test (IAT) in psychology to design prompts that reveal latent biases in LLMs. Rather than directly asking models about their views on particular social groups – which, like humans, they often avoid answering – they instead present structured tasks that require associations or comparative decisions. For instance, a model might be asked to assign positive or negative words to names with racial or gender connotations, or to choose which of two candidates is more suitable for a specific role. This prompt-based approach bypasses explicit self-reports and mirrors the logic of involuntary responses in psychological testing, allowing underlying implicit stereotypes to surface indirectly through model behavior.
@@
Science, Innovation, and LLM-as-a-Judge: Prior work has increasingly applied LLMs to assist researchers by generating feedback, reviews, or reflections throughout the research process [70, 25]. Most recent studies have focused on addressing challenges posed by the rapid growth of academic publishing, such as rising review workloads and declining review quality [5, 51]. Liang et al. [37] evaluate GPT-4’s capacity to generate scientific feedback and find substantial overlap with human reviews, especially in early manuscript stages. Thakkar et al. [60] developed a "Review Feedback Agent" combining multiple LLMs; in a randomized trial, 27% of reviewers revised their feedback after receiving LLM suggestions, which were more detailed and informative. Tyser et al. [62] propose a broader framework, training LLMs to align with human preferences through pairwise comparisons and enhancing review robustness with techniques like adaptive prompting and role-playing. While prior studies on LLMs in peer review have primarily focused on augmenting or replicating human judgment to address scalability and quality concerns, our argument departs from this substitutional framing. Instead of merely simulating reviewers to reduce workload or enhance coverage of aspects, we leverage LLMs’ generative capacity to surface the implicit heuristics and tacit values that often shape peer review decisions but remain unspoken. By prompting LLMs to hypothesize why one paper is rated more highly than another, we reveal not only individual reviewer preferences but also broader cultural norms and evaluative codes that govern scientific judgment. In doing so, we treat AI not as a replacement for human thinking but as a diagnostic tool – one that can reflect, exaggerate, and ultimately make visible the "unwritten code" underpinning peer review and, more broadly, other sociotechnical systems.

### v5

@@ 5 Discussions
It is important to emphasize that our goal is not to advocate for the automation of judgment tasks by LLMs. Rather, our primary goal is to present a position through a conceptual framework — namely, to exploit the biases and preference rationale embedded in LLMs by human discursive culture, then self-amplifying, reproducing, and explicitly identifying who might be harmed by these biases. Much of the existing research has focused on more explicit, easily identifiable forms of biases associated with social dimensions such as gender and race (Ntoutsi et al., 2020). Researchers have developed methods to bypass the barriers when the model refuses to disclose its biases, such as using associated words (Bai et al., 2025). However, they constitute the “prior” itself. These biases and heuristics still manifest superficially and can be uncovered through direct or indirect cues. In contrast, what we focus on in this paper is the "posterior" and how it diverges from the prior. Surprisingly, the results show that even when an LLM’s prior (i.e., its stereotypes) is "good", its posterior can still be "bad" - let alone when its prior is already "bad". When acting as reviewers, LLMs appear to draw on two distinct systems to explain differences in quality signals. One is more normative and less confident in the evaluation process. The other is more narrative-driven, context-sensitive, and ultimately more confident and consistent. When judgment can be achieved through both systems, LLMs are more likely to apply the second system across the board as heuristics because they more consistently explain quality differentials. Humans, like LLMs, may also favor contextual criteria in evaluation, but they articulate this shift in terms of normative standards.
@@
How to articulate the unwritten code in other high-stakes evaluative settings? In hiring, LLMs can be asked to rank resumes: while initial judgments may rely on normative criteria such as education, experience, and skills, prompting models to articulate deeper rationales can surface less explicit but influential factors like perceived cultural fit—for instance, inferring misalignment when a candidate’s prior employers differ culturally from the hiring firm (Rivera, 2012) and being rejected silently. Similarly, in college admissions, encouraging LLMs to explain their evaluations of personal statements enables systematic analysis of implicit reasoning; prior work shows that essays from certain ethnic or linguistic backgrounds exhibit stylistic differences that may be inadvertently penalized (Lee et al., 2025; Alvero et al., 2021; Alvero and Pattichis, 2024). Examining LLM scoring rationales in these contexts can thus help identify how implicit signals contribute to bias in hiring, admissions, and criminal justice.
@@
To what extent does an LLM’s shift from prior to posterior actually reflect human reasoning? Humans may never explicitly confess to using narrative or contextual heuristics in reviewing and other social dimensions, yet we have fairly strong evidence that they do. Prior survey-based studies show that reviewers in the social sciences place substantially greater emphasis on “interpretive challenges that question a study’s framing” than on methodological rigor (Teplitskiy, 2016; Strang and Siler, 2015). We push LLMs to articulate such emphasis and generalize this insight to conference papers in technical fields, where - owing to normative expectations around methodological rigor and SOTA performance - reviewers are less inclined to make such concerns explicit  [^4] https://hackingsemantics.xyz/2020/reviewing-models/ but use "lazy" words as a mask precisely in NLP (Purkayastha et al., 2025). Moreover, if human score variance could be understood as comprising both explicit and implicit components, with LLMs’ prior-to-posterior shift accounting for nearly all variance (as shown in Section 3 "iterative search") while human-stated explicit judgments map only to the prior, then the posterior can reasonably be taken to capture the implicit part.
@@
Another view holds that even if we acknowledge the existence of unwritten codes, making them explicit serves little purpose and may even enable people to “game the system.” But this view neglects the crucial balance between transparency and normativity. The fear that publicizing unwritten codes will be exploited implicitly assumes that it is preferable for such rules to continue operating in the shadows. In reality, the opposite is true: "hidden curriculum" benefits those who already occupy advantaged positions or happen to be in the know (Hariharan, 2019). For example, researchers from elite institutions who are familiar with Western academic norms naturally understand how to write and frame a compelling “story” (Chen et al., 2024b) - they have long benefited from tacit rules without needing them to be made explicit. The purpose of transparency is not to provide a cheat sheet, but to allow marginalized groups to understand the hidden barriers they face and thereby push institutions to reflect on and revise those standards. If people do game the system after unwritten codes are revealed, it is not a failure of transparency — it is evidence that those codes themselves need institutional correction.
@@
Can LLMs replace the peer review and other judge processes? Our goal is not to provide a definitive answer – although certainly not yet. Nevertheless, our work suggests that insofar as LLMs function as cultural technologies that encode societal norms (Blodgett et al., 2020; Farrell et al., 2025), then their "cognitive shift" in peer review reflects analogous unacknowledged shifts in human reviewers. That said, while LLM-generated reviews have been proven useful and cover broader and more specific aspects than human reviews (Liang et al., 2024; Yuan et al., 2022; Thakkar et al., 2025) – likely including those humans may be unwilling or unable to disclose – LLM reviewers still lack constructive feedback essential for critique when judging a standalone paper. Pairing papers for comparative judgment of feature existence and magnitude in some sense addresses this concern, as demonstrated here and in previous experiments (Bai et al., 2025).
What should we do at the implementation level? LLMs cannot be made fully unbiased, but they can be designed to surface, label, and contrast the implicit heuristics they rely on. At an implementation level, this means shifting from single-shot judgments to procedures that explicitly elicit multiple rationales, track prior–posterior shifts, and expose which criteria become dominant only after iterative comparison. Practically, LLMs should be used as diagnostic tools rather than automated judges: systems can require models to generate and evaluate multiple competing hypotheses, retain confidence signals, and highlight systematic gaps between what is explicitly articulated and what is implicitly rewarded. These signals should be logged and presented to human decision-makers, not collapsed into final scores. We should establish evaluation metrics for transparency, not just accuracy. Implementation should emphasize perspective diversity. Using multiple models, prompts, or reviewer roles (e.g., complementary–neutral–proximal perspectives) helps prevent any single normative frame from dominating. Diversity here should be enforced not only at the level of prompts, but at the representational level - through steering, architectural tweaks, or training differences that lead to meaningfully different internal representations and behaviors. Such collective calibration mechanisms allow institutions to recognize, debate, and potentially revise their unwritten codes.
@@
This work complies with the ACL Code of Ethics. We used AI assistants to expedite the coding process. All code snippets produced by AI assistants were verified by the first author before they were incorporated. For writing, we only used AI assistants to check grammar. No external ethics review was required, as the experiments solely analyzed scientific papers, thereby posing no potential harm to individuals across any demographic or geographical group. Our study uses LLMs to surface implicit biases and unwritten evaluative codes in peer review. While this promotes transparency and fairness, it may also risk reinforcing stereotypes if misused. We mitigate this by anonymizing data, framing the method as diagnostic rather than prescriptive, and highlighting the importance of responsible use. All experiments were conducted using two NVIDIA 40-GB A100 GPUs.
@@ Appendix B Human annotation experiments
Here, we complement the labeling methods (LLM labeling (major) and the dictionary method (complementary)) with human annotation. Three master’s students specializing in social computing were involved to label the reviewer comments using exactly the same prompt as GPT-4o (if reviewers’ attitudes toward the aspects in a given hypothesis were positive = 1 or negative = –1, they were considered as “mentioning”, or 0 - not mentioned at all). The annotators are from Asia: two males and one female. They labeled independently, without using AI assistance and without communicating with one another. We find that one person shows high consistency with LLM labels, with an overlap of 0.76. Their majority-voted label across annotators shows an overlap with LLM labels of 0.70. These results show the reliability of LLMs in annotating text data.
@@ Appendix C Results using other LLMs are consistent
To increase the generalizability of our work, we explored other popular open-sourced or closed models, including Deepseek-chat-v3-0324  [^5] https://api-docs.deepseek.com/news/news250325, Mistral 24b instruct  [^6] https://huggingface.co/mistralai/Mistral-Small-24B-Instruct-2501, and Gemini 2.5 Flash-Lite  [^7] https://docs.cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-5-flash-lite. The result of Gemini 2.5 Flash is very similar to that of its lite version., to replicate the full set of experiments. Overall, we find our conclusions to be robust. Researchers have observed increasing convergence among LLMs, largely because major model families are trained on similar mixtures of large-scale web corpora, code repositories, and curated texts (Wu et al., 2025; Wolfram and Schein, 2025). If several LLMs from different companies independently articulate the same belief-change dynamics, this convergence itself may be informative for the biases learned from in-the-wild, human-generated corpora. Specifically:
  • Alignment between priors and human comments, but not posteriors: All three additional models show a relatively consistent set of priors as OpenAI models used in the main paper, and align with human comments about the definition of “good science”, compared to the low alignment for posteriors VS human comments. Human reviewers tend to explicitly reward aspects that moderately align with LLMs’ normative priors: OpenAI correlation = 0.49; Deepseek 0.28; Gemini 0.65; Mistral 0.45. By contrast, human reviewers generally avoid articulating posteriors in their review comments (OpenAI correlation = - 0.14; Deepseek = –0.16; Gemini = -0.20; Mistral = -0.22), but all models do articulate these posteriors.
  • Shift from prior to posterior: During the transition from prior to posterior, we still observe a systematic shift from rigor-related aspects toward storytelling perspectives. All models demonstrate surprisingly high consistency in using certain heuristics such as “whether the paper tells a good interdisciplinary story”, and across models, they show on average 89% agreement with Open-AI’s judgment. This hypothesis also predictably gained the highest prior-to-posterior attention gain: interdisciplinary integration (OpenAI top-1; DeepSeek top-1; Gemini top-2; Mistral top-5) and contextualization (DeepSeek top-2; OpenAI top-2; Gemini top-1; Mistral top-1) are both systematically overestimated in the posterior belief shift. By contrast, rigor is consistently underestimated (OpenAI lowest top-1; DeepSeek lowest top-1; Gemini lowest top-2; Mistral lowest top-2).
@@ Appendix E Fidelity experiments
Summarization using methods from (Zhang et al., 2025) demonstrates high fidelity compared to the raw text. In Figure 7, panel (a), we use the embedding model from Singh et al. (2023), specifically pretrained on scientific papers’ citation relations using contrastive learning. We embed both the summarizations and their corresponding raw texts in each section. To visualize and compare the distribution of these embedding areas, we reduce their dimensionality to two. We then estimate the boundary of each region using a 5% Gaussian kernel density estimate, drawing contours that enclose the densest 95% of data points in each embedding set. If the summarizations contain significant hallucinations, we would expect their embeddings to cover meaningfully different regions than those of the raw texts. However, in panel (a), we find that the distributions largely overlap, suggesting strong semantic alignment. In panel (b), we compute the cosine similarity between the two (raw, summarized) embeddings and find that most values are above 0.95.
@@
We conducted a complementary robustness check to assess whether LLMs make consistent judgments when directly reading full texts, compared to the extended summarization used in the paper for producing generalizable hypotheses under the limited context window. If using full texts from parsed PDFs, the results are essentially identical to those obtained with extended summaries (given their high fidelity), yielding a consistency score of 0.89 (i.e., the same judgment) across 1,000 judgments.
@@ Appendix G Related work
Hypothesis Generation: Recent studies have explored the capabilities of LLMs in generating novel, testable, and interpretable hypotheses across a range of domains. Zhou, Liu et al. (Liu et al., 2024a, 2025; Zhou et al., 2024) proposed a framework that prompts LLMs to generate hypotheses from labeled datasets, beginning with a few examples. These initial hypotheses are iteratively refined using a reward mechanism inspired by the upper confidence bound algorithm from multi-armed bandits, balancing exploration and exploitation. Other works have integrated external knowledge — such as scientific literature, knowledge graphs, and raw web corpus — into the hypothesis generation process (Xiong et al., 2024; Kulkarni et al., 2025; Wang et al., 2023, 2024; Yang et al., 2024), improving the accuracy and novelty of LLM-generated hypotheses while reducing hallucinations. This line of work demonstrates the potential of LLMs in advancing scientific discovery. Beyond computer science, other disciplines are also beginning to explore these possibilities. Ludwig and Mullainathan (Ludwig and Mullainathan, 2024) highlighted the promise of LLMs as tools for hypothesis generation in economics, arguing that machine learning models can uncover patterns and relationships that traditional econometric methods may miss. Manning, Zhu, and Horton (Manning et al., 2024) presented a simulation-based approach for automatically generating and testing social science hypotheses in silico. In each simulated agent-interaction case, causal relationships are both proposed and tested by the LLM system, with results aligning well with those derived from economic theory, particularly in predicting the signs of estimated effects. In medicine, LLMs have been used to hypothesize synergistic drug combinations for breast cancer treatment, with several predictions later validated in laboratory experiments (Bazgir et al., 2025), showcasing the potential of LLMs to surface effective, previously overlooked therapeutic strategies.
@@
Implicit Bias in LLMs: Concerns about bias in AI systems have long focused on the ways these technologies can absorb and amplify harmful content from large-scale human-generated datasets – particularly prejudices related to race, gender, and other social dimensions (Ntoutsi et al., 2020). Large language models (LLMs), trained on vast corpora of human text, are especially prone to perpetuating and even intensifying these biases (Weidinger et al., 2021; Gallegos et al., 2024). Such biases are not computational artifacts but reflect the deeper cultural and social frameworks that shape human communication (Babaeianjelodar et al., 2020; Lee et al., 2024a; Guilbeault et al., 2024). The recognition that LLMs encapsulate and propagate societal biases and heuristics is now well-established in literature, framing these models as inherently cultural and social technologies (Blodgett et al., 2020; Farrell et al., 2025; Weidinger et al., 2021; Naous et al., 2023; Li et al., 2024a). Much of the existing research on LLM bias has concentrated on explicit, easily identifiable forms of bias – such as derogatory language, disparate system performance, erasure, exclusionary norms, misrepresentation, stereotyping, and toxicity (Gallegos et al., 2024; Czarnowska et al., 2021). Meanwhile, efforts in AI alignment, including outer alignment, inner alignment, and interpretability, have sought technological strategies for migrating such risks (Shen et al., 2023; Meade et al., 2021; Kumar et al., 2022). Despite progress in addressing explicit bias, unwritten codes in LLMs’ "cognitive" process remain understudied. Like human cognition, LLMs may exhibit subtle, unconscious forms of bias even when explicit prejudices are absent (Shen et al., 2023; Lin and Li, 2025). These implicit biases are often harder to detect yet can have deeper and more pervasive effects on model outputs and downstream decision-making. Thus, systematically identifying these “unwritten codes" in LLMs is crucial to ensuring fairness, accuracy, and trustworthiness (Lin and Li, 2025). One method worth mentioning in this topic is (Bai et al., 2025). The authors draw on experimental paradigms from the Implicit Association Test (IAT) in psychology to design prompts that reveal latent biases in LLMs. Rather than directly asking models about their views on particular social groups – which, like humans, they often avoid answering – they instead present structured tasks that require associations or comparative decisions. For instance, a model might be asked to assign positive or negative words to names with racial or gender connotations, or to choose which of two candidates is more suitable for a specific role. This prompt-based approach bypasses explicit self-reports and mirrors the logic of involuntary responses in psychological testing, allowing underlying implicit stereotypes to surface indirectly through model behavior.
@@
Science, Innovation, and LLM-as-a-Judge: Prior work has increasingly applied LLMs to assist researchers by generating feedback, reviews, or reflections throughout the research process (Yuksekgonul et al., 2024; Janssen et al., 2023). Most recent studies have focused on addressing challenges posed by the rapid growth of academic publishing, such as rising review workloads and declining review quality (Azad and Banu, 2024; Rogers et al., 2023). Liang et al. (Liang et al., 2024) evaluate GPT-4’s capacity to generate scientific feedback and find substantial overlap with human reviews, especially in early manuscript stages. Thakkar et al. (Thakkar et al., 2025) developed a "Review Feedback Agent" combining multiple LLMs; in a randomized trial, 27% of reviewers revised their feedback after receiving LLM suggestions, which were more detailed and informative. Tyser et al. (Tyser et al., 2024) propose a broader framework, training LLMs to align with human preferences through pairwise comparisons and enhancing review robustness with techniques like adaptive prompting and role-playing. While prior studies on LLMs in peer review have primarily focused on augmenting or replicating human judgment to address scalability and quality concerns, our argument departs from this substitutional framing. Instead of merely simulating reviewers to reduce workload or enhance coverage of aspects, we leverage LLMs’ generative capacity to surface the implicit heuristics and tacit values that often shape peer review decisions but remain unspoken. By prompting LLMs to hypothesize why one paper is rated more highly than another, we reveal not only individual reviewer preferences but also broader cultural norms and evaluative codes that govern scientific judgment. In doing so, we treat AI not as a replacement for human thinking but as a diagnostic tool – one that can reflect, exaggerate, and ultimately make visible the "unwritten code" underpinning peer review and, more broadly, other sociotechnical systems.